OpenStudy (anonymous):

Solve integral:

OpenStudy (anonymous):

\[\int\limits x/(x^3-1)dx\]

OpenStudy (anonymous):

you can replace x with sin x then x= sinx and \[x ^{2} -1 =\sin ^{2} -1 = \cos ^{2}\] and dx=cos x dx then the integration becomes \[\int\limits_{}^{}sinx /\cos ^{2}x \times \cos x dx = \int\limits_{}^{}\tan x dx\]

OpenStudy (anonymous):

i think its x^3

OpenStudy (anonymous):

@El-naggar

OpenStudy (anonymous):

yes x^3

OpenStudy (anonymous):

i think better to use partial fractions....not sure....

OpenStudy (anonymous):

sorry i saw it x^2

OpenStudy (anonymous):

@El-naggar : it's okay

OpenStudy (anonymous):

I got it. Gimmie 5 mins..

OpenStudy (anonymous):

First split the equation into two, as thus: \[{x \over (x-1)(x^2+x+1)} = {A \over (x-1)} + {Bx + C \over (x^2 + x + 1}\] We get that \[A(x^2 + x + 1) + (x-1)(Bx + C) = x\] A = C, A= -B, B= -1/3 \[{x \over (x-1)(x^2+x+1)} = {1 \over 3}{1 \over (x-1)} + {1 \over 3}{-x + 1 \over (x^2 + x + 1)}\] The integrals are now trivial, theyre both log!

OpenStudy (anonymous):

oh i got it Thank again @soati :)

OpenStudy (anonymous):

:) always a pleasure

OpenStudy (anonymous):

:D